MEMORANDUM May 19, 1999
TO: R. Fisher, J. Ford, F. Ghigo, R. Lacasse, R. Norrod, M. McKinnon, C. Niday, T. Weadon
FROM: R. Lacasse
SUBJ.: Minutes of Interference Protection Group 05/18/99 Meeting
In attendance: Fisher, Ford, Lacasse, Niday, Weadon
Minutes of Previous Meeting
Call Outs for Interference
The 140' experienced some intermittent RFI at 8695 MHz. The source was not found; the RFI went away before we could take time to do an azimuth scan.
Future call-outs for interference will be handled by Fisher with Lacasse and Niday as alternates. Both Sizemore and Niday are experienced in navigating the FCC's database to determine frequency assignments.
Action Item Review
1 - Do we need formal coordination of observing with known transmissions on site (OVLBI, antenna range, etc.)?
Several suggestions came about as a result of discussing this question. First, since we are primarily a receiving site, the onus of notification should be on those who are transmitting. This could be easily accomplished with an e-mail exploder. Also, it would not be difficult to have the receive frequencies for all antennas posted on a web page, and updated in real time by the Monitor and Control System. Thirdly, anyone transmitting should check into whether they need a license. No one was aware of whether the Interferometer or the 20-Meter had significant problems with interference. Ghigo needs to be consulted on this.
2 - Measurement of OVLBI transmission field strenghts at the GBT.
Phil Jewell has requested that we measure transmission field strengths from the OVLBI antenna. These would serve as useful guides in coordinating OVLBI and GBT observing. Fisher will get the transmission parameters from Minter. Niday will round up equipment. All will help with measurements.
3 - Progress on RFI Suppression from equipment that has been measured in the anechoic chamber
Several iterations of the new emergency stop system have been tested. The final version has been tested by Pedtke and was found satisfactory. Niday will verify this test.
Parts of the laser ranging system will be tested after today's meeting.
Video cameras have been tested recently, in a modified enclosure and are clean when powered by batteries inside the enclosure. Cables need to be brought out and further testing conducted.
The IF rack is undergoing a complete rebuild with RFI suppression in mind. No tests have been conducted yet.
Test results of other equipment destined for the equipment room need to be reviewed in light of the tests that were conducted on the receiver building.
4 - Site Radio System
Operating frequencies have been found. Equipment assembly has begun. A transmission test from the water tower is being coordinated with Holstine. Harmonics need to be tested. A resistive combiner is being constructed to combine the outputs of several transmitters for transmission by one antenna. The system should be operational this summer.
5 - Monitor station and anechoic chamber development
New poles for the antennas for the monitor station will be installed shortly. Cables have been ordered. The overall plan is to have a separate receiver and LO for each octave.
To test for FCC compliance, a turn-table is required in the anechoic chamber. Egan will look into mechanical details. Weadon suggested the use of a pneumatic motor to minimize interference.
Weadon and Lacasse need to test the continuum detector system ASAP.
6 - Microwave Oven Enclosure
A commercial enclosure was purchased ($1500) from EMCOR. It was received recently and found to be poorly constructed. It will be returned to the vendor for repair. Ford has since examined equipment from AMCO and found it similar in price but superior in quality.
7 - Shielded Room Entrances
Some absorber has been found, but not enough to build a prototype. It is possible to build a prototype at K-band and then scale the design to be effective at lower frequencies.
9 - Use of Interference Reporting System
Fisher will check to see if interference reporting system was used to document interference at 8695 MHz referred to above.
Agenda items 8, and 9 to 15 were not discussed.
Next Meeting: June 1, 0800.